# FusionBench: A Comprehensive Benchmark of Deep Model Fusion

### v3

@@ Appendix C Generalization Experiments
  • In the second set of experiments, we merge models trained six tasks (SUN397, Cars, GTSRB, EuroSAT, DTD, MNIST) and evaluate the fused model on the unseen tasks (RESISC45, SVHN). The results are shown in Table 16.
By conducting these two sets of generalization experiments, we aim to gain a comprehensive understanding of how the CLIP-ViT-B/32 models, when fused with knowledge from different task combinations, can perform on various unseen tasks. From these experimental results, we can observe instances of negative transfer when evaluating the fused CLIP models on unseen tasks. Here, negative transfer occurs when the knowledge gained from fine-tuning on a set of tasks hinders the model’s performance on new, unseen tasks. In other words, the model’s ability to generalize and adapt to novel challenges is compromised due to the specific knowledge acquired during the fine-tuning process. The presence of negative transfer in these experiments highlights the challenges and limitations of model fusion and generalization in the context of the CLIP-ViT-B/32 models. Several factors can contribute to negative transfer, such as:
  • Task dissimilarity: If the unseen tasks are significantly different from the tasks used for fine-tuning, the learned representations may not be directly applicable, leading to performance degradation.
  • Overspecialization: Fine-tuning on a specific set of tasks may cause the model to overfit to task-specific features and patterns, reducing its ability to generalize to new tasks.
  • Interference between tasks: When merging knowledge from multiple tasks, there may be conflicts or interference between the learned representations, hindering the model’s ability to adapt to unseen tasks effectively.
To mitigate the negative transfer and improve the generalization ability of merged models, several strategies can be explored, such as:
  • Task selection: Carefully selecting tasks that are more similar or complementary to the target unseen tasks can help reduce the risk of negative transfer. This is adapted in Wu et al. [58], where the Fisher information matrix is computed for a proxy metric for task similarity.
  • Regularization techniques: Applying regularization methods, such as weight decay or dropout, during the fine-tuning process may help prevent overfitting and promote better generalization.
Appendix D Prompt-Based Text-to-Text Generation
@@
This section details the prompt templates employed for each of the eight text-to-text generation tasks from the GLUE benchmark., see Section 3.3 for more details. Within each task, we provide the format of the input text, and the corresponding target text mapping. These templates are crucial in fine-tuning the Flan-T5 models for generating appropriate text outputs tailored to each specific task.

### v4

@@ E.4.1 Generalization Experiments
  • In the second set of experiments, we merge models trained on six tasks (SUN397, Cars, GTSRB, EuroSAT, DTD, MNIST) and evaluate the fused model on the unseen tasks (RESISC45, SVHN). The results are shown in Table 15.
Tables 14 and 15 present the generalization performance of various multi-task model fusion algorithms when merging CLIP-ViT-B/32 models trained on six seen tasks and evaluating their performance on two unseen tasks. This analysis helps us understand how well the fused models can adapt to new tasks that were not encountered during the training and model fusion process.
E.4.2 Robustness Experiments
Table 16: Ablations of the test data distribution on ViT-B/32 (for all methods, $\lambda=0.3$). Here underlined method names indicate methods that require test-time adaptation training, while italicized method names indicate model mixing approaches.
Clean Test Set Corrupted Test Set (Motion Blur)
Fisher Merging 66.0 92.7 83.7 78.7 80.3 60.7 57.6 81.7 78.4 69.6
RegMean 72.1 97.5 88.9 93.9 88.1 70.0 71.3 87.5 86.8 78.9
Task Arithmetic 64.6 91.8 80.2 74.8 77.9 62.4 59.2 78.5 63.3 65.9
Ties-Merging 65.2 83.3 78.1 67.4 73.5 64.4 53.9 76.4 57.1 62.9
AdaMerging 75.2 94.3 87.6 96.7 88.5 72.4 72.7 85.3 94.3 81.2
WEMoE 77.4 98.9 94.4 99.0 92.4 76.5 74.2 93.7 97.4 85.5
Fisher Merging 61.5 50.0 74.7 52.6 59.7 61.6 48.1 76.0 51.3 59.3
RegMean 66.9 51.0 80.6 68.7 66.8 69.4 41.8 84.0 67.7 65.7
Task Arithmetic 59.8 53.3 72.3 45.0 57.6 61.5 52.5 75.0 50.1 59.8
Ties-Merging 60.2 45.6 69.8 38.3 53.5 61.8 47.3 73.1 42.3 56.1
AdaMerging 69.2 40.0 79.6 83.3 68.0 70.0 53.3 82.1 80.0 71.4
WEMoE 75.1 9.7 91.5 91.8 67.0 76.5 9.6 92.7 88.7 66.8
Corrupted Test Set (Pixelate) Corrupted Test Set (Spatter)
Fisher Merging 2.2 34.0 17.0 63.2 29.1 61.4 64.2 74.6 47.3 61.9
RegMean 2.3 38.3 18.2 89.4 37.0 67.7 60.0 81.3 81.9 72.7
Task Arithmetic 2.3 33.2 19.1 65.6 30.0 61.0 62.5 72.8 57.0 63.3
Ties-Merging 3.3 31.8 18.0 58.5 27.9 61.3 52.9 70.3 48.1 58.2
AdaMerging 1.3 52.9 21.0 91.0 41.5 68.4 55.9 78.3 92.3 73.7
WEMoE 0.5 11.6 2.3 97.5 28.0 75.1 9.7 91.4 96.3 68.1
Corrupted Test Set (Contrast) Corrupted Test Set (JPEG Compression)
Fisher Merging 63.8 58.4 75.5 70.4 67.0 66.3 67.6 82.6 58.9 68.8
RegMean 69.6 64.8 84.4 90.0 77.2 71.5 72.6 88.7 82.2 78.7
Task Arithmetic 62.3 55.7 75.3 70.8 66.0 63.9 66.1 80.1 61.0 67.8
Ties-Merging 64.2 52.4 74.8 63.5 63.7 65.0 59.5 77.9 53.2 63.9
AdaMerging 73.1 67.4 83.0 96.2 79.9 72.9 70.7 86.3 90.6 80.1
WEMoE 77.2 34.7 93.1 98.4 75.9 77.3 61.0 94.1 95.7 82.0
To assess the robustness and real-world applicability of multi-task model fusion algorithms, we conduct evaluations on corrupted test sets as presented in Table 16. These experiments are crucial for understanding how fusion methods perform under distribution shifts and data quality degradation, which are common challenges in practical deployment scenarios. We evaluate six different corruption types across four representative image classification tasks (Cars, EuroSAT, RESISC45, GTSRB), including motion blur, impulse noise, Gaussian noise, pixelation, spatter, and JPEG compression.
The results reveal varying degrees of robustness across different fusion methods, with some approaches maintaining reasonable performance under mild corruptions while experiencing significant degradation under severe distortions such as pixelation. Notably, test-time adaptation-based methods like AdaMerging and WEMoE demonstrate better resilience compared to simpler averaging techniques, though all methods show substantial performance drops under extreme corruption conditions, highlighting the ongoing challenge of maintaining model robustness in multi-task fusion scenarios.
E.5 Scaling to Large-Scale Neural Networks
Table 17: Comparison of individual Mistral-7B models and the upscaled model on various benchmark tasks. For our method, we set $k_{gate}=8,k=512$, and the total parameter count is 11.2B. For a better comparison, we also include the performance of the Qwen1.5-14B model as a reference.
Model MMLU TruthfulQA GSM8K ARC Challenge
Mistral-7B-v0.1 (pre-trained) 59.64 42.62 38.81 53.92
Qwen1.5-14B (reference) 66.11 52.00 69.37 49.93
MetaMath-Mistral-7B 60.56 44.79 71.49 51.02
dolphin-2.1-mistral-7b 60.56 55.88 56.93 57.00
speechless-code-mistral-7b-v1.0 61.18 47.47 48.98 57.68
Simple Average 61.42 49.95 67.40 57.59
Task Arithmetic ($\lambda=0.3$) 61.29 49.38 66.94 57.94
SMILE Upscaled model ($k=512$, 11.2B) 60.66 52.79 67.85 54.35
SMILE Upscaled model (Dense experts) 60.61 54.23 70.66 55.12
Model fusion methods can also be applied to large-scale neural networks including Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs). The high computational cost associated with developing LLMs are a significant practical challenge for many researchers. In FusionBench, we have developed multiple model fusion techniques applicable to LLMs for cheap and efficient model scaling, which can save computational resources for developing new models. Moreover, FusionBench is actively developing and integrating more LLM and MLLM merging methods, with ongoing work to expand benchmark coverage to large-scale model fusion tasks. While evaluation methodologies for LLMs and MLLMs are still maturing and computational costs remain a challenge, FusionBench supports established evaluation frameworks such as LM-Evaluation-Harness (eval-harness) for assessing model performance. As the toolkit evolves, more comprehensive LLM task evaluations will be included in future releases.
Take merging Mistral-7B models using SMILE upscaling (tang2024smile) as an example, we compare the performance of the individual Mistral-7B models and the upscaled model on various benchmark tasks in Table 17, as well as the performance of the Qwen1.5-14B model as a reference. The LM-Evaluation-Harness is utilized to assess the performance of the models. We fuse three Mistral-7B models, each fine-tuned for a distinct downstream task as showing varying performance across different tasks, thereby incorporating task-specific expertise. It is observed that the upscaled models and merged methods (Simple Average, Task Arithmetic) generally enhance performance compared to individual models, demonstrating the benefits of model fusion techniques.
Appendix F Prompt-Based Text-to-Text Generation
@@
This section details the prompt templates employed for each of the eight text-to-text generation tasks from the GLUE benchmark, see Section E.3 for more details. Within each task, we provide the format of the input text, and the corresponding target text mapping. These templates are crucial in fine-tuning the Flan-T5 models for generating appropriate text outputs tailored to each specific task.
@@
Reporting Metrics:
We report accuracy for all tasks except for STSB, where we use Spearman’s $\rho$ as the evaluation metric. For task STSB, the model is expected to output a numerical value. An example from the STSB task is as follows:
  • Input:
    • Sentence 1: A plane is taking off.
    • Sentence 2: An air plane is taking off.
  • Output:
    • label: 5
We try to parse the output as a numerical value. If the model outputs a numerical value, we can calculate the Spearman’s rho between the predicted numerical value and the ground truth numerical value. If the model outputs a non-numerical value, we assume the Spearman’s rho is 0, indicating that there is no discernible monotonic increasing or decreasing relationship between the model’s predictions and the ground truth. This is a conservative approach, as even non-numerical outputs might contain some relevant information that’s being discarded in this evaluation.